OpenStudy (anonymous):

Write an equation of a line perpendicular to y = 1/2x - 5 in slope-intercept form that passes through the point (2, -7). (5 points) Select one: a. y=1/2 x - 7 b.y= 2x + 2 c. y= -2x - 5 d. y= -2x - 3

OpenStudy (anonymous):

@Nurali @g

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Nurali can you pleashelp me with a couple other also please ill medal

OpenStudy (nurali):

y = mx + c where m = slope y = (1/2)x - 5 m1 = 1/2 For perpendicular lines, m2 = -1/m1 where m2 = slope for line 2(perpendicular line) m2 = -1/(1/2) = -2 y = -2x + c If the line passes through (2,-7), When x = 2, y = -7 -7 = -2(2) + c -7 = -4 + c c = - 3; Thus, y = -2x - 3

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

My Pleasure.